# Spatial-temporal Memories Enhanced Graph Autoencoder for Anomaly Detection in Dynamic Graphs

### v1

@@ IV-B Spatial-Temporal Memory Learning
The spatial-temporal memory learning aims to capture and store prototypical spatial patterns and temporal patterns of normal node embeddings. Both spatial and temporal memory banks contain $p$ memory items of dimension $D^{\prime}$, denoted as $\mathbf{M}_{sp}\in\mathbb{R}^{P_{s}\times D^{\prime}}$ and $\mathbf{M}_{tp}\in\mathbb{R}^{P_{t}\times D^{\prime}}$, respectively. $\mathbf{M}_{sp}$ and $\mathbf{M}_{tp}$ share the same memory update and read procedures. For simplicity, we only present the learning process for $\mathbf{M}_{tp}$.
IV-B1 Memory Read
We denote $\mathbf{m}_{p}\in\mathbb{R}^{D^{\prime}}(p=1,\ldots,P_{t})$ as the item in the memory $\mathbf{M}_{tp}$, and $\mathbf{z}_{i}\in\mathbb{R}^{D^{\prime}}(i=1,\dots,\tau^{\prime}\times N_{t})$ as the item in encoded node features. As shown in Fig. 3(a), the reading process begins by calculating the attention weights between each node feature $\mathbf{z}_{i}$ and all memory items $\mathbf{m}_{p}$. Prior research [34, 35] primarily adopts cosine similarity to compute self-attention, which restricts the capability to explore the relations between node features and diverse spatial and temporal memory items. To address this problem, we employ a mutual attention mechanism:
$\mathbf{k}_{p}$ $=\mathbf{m}_{p}\mathbf{W}_{K},$ (7)
$\mathbf{q}_{i}$ $=\mathbf{z}_{i}\mathbf{W}_{Q},$ (8)
$\mathbf{v}_{i}$ $=\mathbf{z}_{i}\mathbf{W}_{V},$ (9)
@@ IV-B1 Memory Read
$w_{(i,p)}=\frac{\text{exp}\bigl{(}\mathbf{q}_{i}(\mathbf{k}_{p})^{T}\cdot\frac{1}{\sqrt{D^{\prime}}}\bigr{)}}{\sum^{P}_{p^{\prime}=1}\text{exp}\bigl{(}\mathbf{q}_{i}(\mathbf{k}_{p^{\prime}})^{T}\cdot\frac{1}{\sqrt{D^{\prime}}}\bigr{)}}.\$ (10)
For each node feature $\mathbf{z}_{i}$, we read the memory by a weighted average of the items $\mathbf{m}_{p}$ with the corresponding weights $w_{(i,p)}$, and obtain the readout memory item $\hat{\mathbf{m}}^{t}_{i}\in\mathbb{R}^{D^{\prime}}$ as follows:
$\hat{\mathbf{m}}_{i}=\sum_{p=1}^{P}w_{(i,p)}\mathbf{m}_{p},$ (11)
IV-B2 Memory Update
During the training phase, the memory bank will also be updated to record the spatial and temporal prototypes of normal nodes. As shown in Fig. 3(b), for each memory item $\mathbf{m}_{p}$, we select the node features that are nearest to it based on the matching weights $\mu_{(i,p)}$ as follows:
@@ IV-B2 Memory Update
$\mathbf{m}_{p}\leftarrow\mathbf{m}_{p}+\sum_{i=1}^{K}\mu_{(i,p)}\mathbf{v}_{i}.$ (13)
@@
Figure 3: The illustration of memory read and memory update procedures in the temporal memory module.
@@ IV-C Spatial-temporal Graph Decoder
After the memory read procedure in both spatial and temporal memory modules, we obtain the averaged memory matrices $\hat{\mathbf{M}}_{tp}\in\mathbb{R}^{\tau^{\prime}\times N_{t}\times D^{\prime}}$ and $\hat{\mathbf{M}}_{sp}\in\mathbb{R}^{\tau\times N_{t}\times D^{\prime}}$. Each item within these matrices represents the averaged spatial and temporal normal prototypes associated with the corresponding node. In this section, we reconstruct the original graph stream with the memory matrices and the encoded latent representation $\mathbf{Z}$.
As shown in Fig. 2, our method inputs the concatenation of $\hat{\mathbf{M}}_{tp}$ and $\mathbf{Z}$ into the gated temporal decoder, which is the reverse process of equation 4, and outputs $\hat{\mathbf{Z}}\in\mathbb{R}^{\tau\times N_{t}\times D^{\prime}}$. Then we concatenate $\hat{\mathbf{M}}_{sp}$ with $\hat{\mathbf{Z}}$ and input it into the spatial decoder. We use an L-layer GCN as the attributive decoder. The $l$-th layer is formatted as follows:
[…]
where $\parallel$ denotes concatenation operation. $\hat{\mathbf{H}}^{(t,L)}$ is $\hat{\mathbf{X}}^{t}$. We employ the inner product of the concatenation of $\hat{\mathbf{M}}_{sp}$ and $\hat{\mathbf{Z}}$ as the structural decoder, formatted as follows:
$\hat{\mathbf{A}}^{t}=\sigma\bigl{(}\hat{\mathbf{M}}_{sp}[t,:,:]\parallel\hat{\mathbf{Z}}[t,:,:]\bigr{)}\bigl{(}\bigl{(}\hat{\mathbf{M}}_{sp}[t,:,:]\parallel\hat{\mathbf{Z}}[t,:,:]\bigr{)}\bigr{)}^{T}.$ (15)
Where $\sigma(\cdot)$ denotes the Sigmoid activation function.
@@ IV-D Unified Anomaly Detector
During inference, the loss in equation (19) is adopted as anomaly score. The anomaly score for each node is calculated $R$ times to ensure that the final anomaly scores are statistically stable and reliable.
@@ 5             Extract the graph sequence $\{\mathcal{G}^{t-\tau+1},\ldots,\mathcal{G}^{t}\}$;
21             Calculate the anomaly score for each node $v_{i}$ via Eq. (19).
@@ IV-E Complexity Analysis
In this subsection, we analyze the time complexity of each component in the STRIPE framework. We employ an L-layer GCN for spatial encoding and decoding of a graph sequence with window size $\tau$, which brings a complexity of $\mathcal{O}(\tau LM_{t}D^{\prime}+\tau LN_{t}D^{\prime 2})$, where $M_{t}$ and $N_{t}$ are the averaged edge and node number for each snapshot. For memory read and update, the complexity is mainly caused by mutual attention mechanism, which is $\mathcal{O}(\tau N_{t}^{2}D^{\prime})$. For gated temporal convolution, the time complexity is $\mathcal{O}(\tau LN_{t}^{2}D^{\prime})$, which is far less than the other components and can be ignored. Therefore, to apply an L-layer STRIPE to a graph sequence of window size $\tau$, the overall time complexity is $\mathcal{O}\bigl{(}\tau L(M_{t}D^{\prime}+N_{t}D^{\prime 2}+N_{t}^{2}D^{\prime})\bigr{)}$.
@@ V Experimental Study
In this section, we conduct extensive experiments on four real-world benchmark datasets to evaluate the performance of STRIPE. Specifically, from sectionV-A to V-D, we introduce the experimental setups. Then in section V-E, we compare our method with the state-of-the-art baseline methods on the node anomaly detection task. After that, we conduct an ablation study to validate the effectiveness of each component of STRIPE in section V-F. In section V-G, we study the parameter sensitivity to further investigate the property of STRIPE. We also demonstrate the effectiveness of the proposed spatial and temporal memory modules with a case study in V-H.
@@ V-A Datasets
We assess the performance of STRIPE and its competitors on four real-world temporal networks, including two co-author networks, a social network, and a biological network. The description of the datasets is shown in Table II.
  • DBLP-3 and DBLP-5111https://dblp.uni-trier.de are co-author networks where the nodes represent the authors and the edges represent the co-authorship between them. The node attributes in a network snapshot are extracted from the titles and abstracts of the corresponding author’s publications during a period by word2vec. The authors in DBLP-3 and DBLP-5 are from three and five research areas, respectively.
  • Reddit [^2] https://www.reddit.com/ is a social network where the nodes represent the posts. Two posts are connected if the same user comments on both posts. Two nodes are connected in a network snapshot if their corresponding posts contain similar keywords during a time period. We apply word2vec to the comments of a post to generate its node attributes
  • Brain is a biological network, with nodes symbolizing distinct cubes of brain tissue and edges reflecting their connectivity. The network is derived from task-based functional magnetic resonance imaging (fMRI) data [^3] https://tinyurl.com/y4hhw8ro collected in real-world settings. This fMRI data is gathered while subjects perform various tasks sequentially. For each network snapshot, we utilize Principal Component Analysis (PCA) on the fMRI data of a specific time period to determine the attributes of each node. Connectivity between two nodes is established if they exhibit a similar level of activation during the observed time frame.
Due to the lack of ground-truth anomalies in these datasets, we need to manually inject synthetic anomalies into the original networks for evaluation in the testing phase. For a fair comparison, we follow the anomaly injection strategies used in [36] and [11], and inject equal numbers of structural anomalies and attributive anomalies for each snapshot $\mathcal{G}^{t}$ in the test set:
  • Structural anomaly injection. Following [36], we generate structural anomalies by randomly selecting $N_{p}$ nodes from node set $\mathcal{V}^{t}$ and connecting them to form fully connected cliques. The selected $N_{p}$ nodes are labeled as structural anomaly nodes. This process is repeated $q$ times to generate $q$ cliques. Following [36], we fix $N_{p}$ as 15 and set $q$ to 7, 11, 14, and 8 for DBLP-3, DBLP-5, Reddit, and Brain, respectively.
  • Attributive anomaly injection. Following [11], attributive anomalies are created by randomly selecting $N_{p}\times q$ nodes from $\mathcal{V}^{t}$. For each chosen node $v^{t}_{i}$, we sample an additional $k$ nodes to form a candidate set: $\mathcal{V}^{t}_{i,attr}=\{v^{t}_{1},\ldots,v^{t}_{k}\}$. Then, we replace the feature vector of $v_{i}$ with the node feature from $\mathcal{V}^{t}_{i,attr}$ that has the largest attribute distance from $v^{t}_{i}$. Following [11], we set $s$=2 and $k$=50 for all the datasets.
@@
TABLE II: Statistics of the Datasets. AR represents the anomaly ratio, calculated as the ratio of the number of anomalies to the total number of nodes.
Datasets Nodes Edges Attributes Timestamps Anomalies AR
DBLP-3 4,257 38,240 100 10 210 0.049
DBLP-5 6,606 65,915 100 10 330 0.050
Reddit 8,291 292,400 20 10 420 0.051
Brain 5,000 1,975,648 20 12 240 0.048

### v2

@@ IV-B Spatial-Temporal Memory Learning
The spatial-temporal memory learning aims to capture and store prototypical spatial patterns and temporal patterns of normal node embeddings. Spatial and temporal memory banks contain $\tau P_{s}$ and $P_{t}$ memory items of dimension $D^{\prime}$, denoted as $\mathbf{M}_{sp}\in\mathbb{R}^{\tau\times P_{s}\times D^{\prime}}$ and $\mathbf{M}_{tp}\in\mathbb{R}^{P_{t}\times D^{\prime}}$, respectively. In the rest of this section, we introduce the spatial and temporal memory modules, respectively.
IV-B1 Spatial Memory Module
We denote $\mathbf{m}^{r}_{p}\in\mathbb{R}^{D^{\prime}}(p=1,\ldots,P_{s};r=t-\tau+1,\ldots,t)$ as the item of memory $\mathbf{M}_{sp}$, and $\mathbf{h}^{r}_{i}\in\mathbb{R}^{D^{\prime}}(i=1,\dots,N_{t};r=t-\tau+1,\ldots,t)$ as the spatial encoded feature of node $i$ at time $r$.
Memory Read.. As shown in Fig. 3(a), the reading process begins by calculating the attention weights between each node feature $\mathbf{h}^{r}_{i}$ and all memory items $\mathbf{m}^{r}_{p}$. Prior research [45, 46] primarily adopts cosine similarity to compute self-attention, which restricts the capability to explore the relations between node features and diverse spatial and temporal memory items. To address this problem, we employ a mutual attention mechanism:
$\mathbf{k}_{p}$ $=\mathbf{m}^{r}_{p}\mathbf{W}_{K},$ (7)
$\mathbf{q}_{i}$ $=\mathbf{h}^{r}_{i}\mathbf{W}_{Q},$ (8)
$\mathbf{v}_{i}$ $=\mathbf{h}^{r}_{i}\mathbf{W}_{V},$ (9)
@@ IV-B1 Spatial Memory Module
$w_{(i,p)}=\frac{\text{exp}\bigl{(}\mathbf{q}_{i}(\mathbf{k}_{p})^{T}\cdot\frac{1}{\sqrt{D^{\prime}}}\bigr{)}}{\sum^{P_{s}}_{p^{\prime}=1}\text{exp}\bigl{(}\mathbf{q}_{i}(\mathbf{k}_{p^{\prime}})^{T}\cdot\frac{1}{\sqrt{D^{\prime}}}\bigr{)}}.\$ (10)
For each node feature $\mathbf{h}^{r}_{i}$, we read the memory by a weighted average of the items $\mathbf{m}^{r}_{p}$ with the corresponding weights $w_{(i,p)}$, and obtain the readout memory item $\hat{\mathbf{m}}^{r}_{i}\in\mathbb{R}^{D^{\prime}}$ as follows:
$\hat{\mathbf{m}}^{r}_{i}=\sum_{p=1}^{P_{s}}w_{(i,p)}\mathbf{m}^{r}_{p},$ (11)
Assigning this procedure to all $i\in[1,N_{t}]$ and $r\in[t-\tau+1,t]$, we obtain the readout memory matrix $\hat{\mathbf{M}}_{sp}\in\mathbb{R}^{\tau\times N_{t}\times D^{\prime}}$.
Memory Update. During the training phase, the memory bank will also be updated to record the spatial prototypes of normal nodes. As shown in Fig. 3(b), for each memory item $\mathbf{m}^{r}_{p}$, we select the node features that are nearest to it based on the matching weights $\mu_{(i,p)}$ as follows:
@@
$\mathbf{m}^{r}_{p}\leftarrow\mathbf{m}^{r}_{p}+\sum_{i=1}^{K}\mu_{(i,p)}\mathbf{v}_{i}.$ (13)
IV-B2 Temporal Memory Module
Unlike spatial memories that capture normal patterns within each graph snapshot, temporal memories aim to characterize the prototypical pattern of evolving trends among different snapshots. Specifically, $m_{q}\in\mathbb{R}^{D^{\prime}}$ is the item of memory bank $\mathbf{M}_{tp}$. For each temporal node feature $\mathbf{z}_{i}\in\mathbb{R}^{D^{\prime}}(i=1,\dots,\tau^{\prime}\times N_{t})$, its corresponding readout memory $\hat{\mathbf{m}}_{i}$ is calculated as follows:
$\hat{\mathbf{m}}_{i}=\sum_{q=1}^{P_{t}}w_{(i,q)}\mathbf{m}_{q}.$ (14)
We adopt the same memory read procedure as in Eq. (10) to calculate $w_{(i,q)}$ and utilize Eq. (12) to update temporal memories. The overall readout matrix for temporal features is denoted as $\hat{\mathbf{M}}_{sp}\in\mathbb{R}^{\tau\times N_{t}\times D^{\prime}}$. Each item within these matrices represents the averaged spatial and temporal normal prototypes associated with the corresponding node.
@@ IV-B2 Temporal Memory Module
Figure 3: The illustration of (a) memory read and (b) memory update procedures in the spatial memory module.
@@ IV-C Spatial-temporal Graph Decoder
In this section, we reconstruct the original graph stream with the averaged memory matrices and latent representation $\mathbf{Z}$. As shown in Fig. 2, our method inputs the concatenation of $\hat{\mathbf{M}}_{tp}$ and $\mathbf{Z}$ into the gated temporal decoder, which is the reverse process of equation (3), and outputs $\hat{\mathbf{Z}}\in\mathbb{R}^{\tau\times N_{t}\times D^{\prime}}$. Then we concatenate $\hat{\mathbf{M}}_{sp}$ with $\hat{\mathbf{Z}}$ and input it into a one-layer MLP to obtain $\{\hat{\mathbf{H}}\}^{t}_{t-\tau+1}$:
$\hat{\mathbf{H}}^{t}=MLP\bigl{(}\hat{\mathbf{M}}_{sp}[t,:,:]\parallel\hat{\mathbf{Z}}[t,:,:]\bigr{)},\$ (15)
where $\parallel$ denotes concatenation operation and $\hat{\mathbf{H}}^{t}\in\mathbb{R}^{N_{t}\times D^{\prime}}$. Then, we use an L-layer GCN as the attributive decoder to reconstruct attribute matrix $\hat{\mathbf{X}}^{t}$:
$\hat{\mathbf{X}}^{t}=\textit{GCN}\Bigl{(}\mathbf{A}^{t},\hat{\mathbf{H}}^{t};\mathbf{\Theta}_{de}^{(l)}\Bigr{)},$ (16)
We employ the inner product of $\hat{\mathbf{H}}^{t}$ as the structural decoder, formatted as follows:
$\hat{\mathbf{A}}^{t}=\sigma\Bigl{(}\hat{\mathbf{H}}^{t}\mathbf{W}_{de}\bigl{(}\hat{\mathbf{H}}^{t}\bigr{)}^{T}\Bigr{)}.$ (17)
Where $\mathbf{W}_{de}\in\mathbb{R}^{D^{\prime}\times D^{\prime}}$ is the weight matrix and $\sigma(\cdot)$ denotes the Sigmoid activation function.
@@ IV-D Unified Anomaly Detector
During inference, the sum of $\mathcal{L}_{rec}$ and $\mathcal{L}_{com}$ is adopted as the anomaly score. The anomaly score for each node is calculated $R$ times to ensure that the final anomaly scores are statistically stable and reliable.
@@ 4             Extract the graph sequence $\{\mathcal{G}^{t-\tau+1},\ldots,\mathcal{G}^{t}\}$;
21             Calculate the anomaly score for each node $v_{i}$ via $\mathcal{L}_{rec}$ and $\mathcal{L}_{com}$.
@@ IV-E Complexity Analysis
In this subsection, we analyze the time complexity of each component in the STRIPE framework. We employ an L-layer GCN for spatial encoding and decoding of a graph sequence with window size $\tau$, which brings a complexity of $\mathcal{O}(\tau LM_{t}D^{\prime}+\tau LN_{t}D^{\prime 2})$, where $M_{t}$ and $N_{t}$ are the averaged edge and node number for each snapshot. For memory read and update, the complexity is mainly caused by the mutual attention mechanism between node features and memory items, which is $\mathcal{O}(\tau N_{t}P_{t}D^{\prime})$. The time complexity for an L-layer gated temporal convolution is $\mathcal{O}(\tau LN_{t}D^{\prime 2})$. Therefore, to apply an L-layer STRIPE to a graph sequence of window size $\tau$, the overall time complexity is $\mathcal{O}\bigl{(}\tau L(M_{t}D^{\prime}+N_{t}D^{\prime 2})+\tau N_{t}P_{t}D^{\prime})\bigr{)}$. As investigated in section V-H2, $P_{t}$ can be restricted to a very limited number, thereby the time complexity of STRIPE is approximately linear to node number. Section V-F provide an empirical analysis of model efficiency.
@@ V Experimental Study
In this section, we conduct extensive experiments on six real-world benchmark datasets to evaluate the performance of STRIPE. Specifically, from sectionV-A to V-C, we introduce the experimental setups. Then in section V-D and V-E, we compare our method with the state-of-the-art baseline methods on node anomaly detection task. We then evaluate time efficiency of the model in section V-F and conduct ablation study to validate the effectiveness of each component of STRIPE in section V-G. In section V-H, we study the parameter sensitivity to further investigate the property of STRIPE. We also demonstrate the effectiveness of the proposed spatial and temporal memory modules with a case study in V-I.
@@ V-A Datasets
We assess the performance of STRIPE and its competitors on six real-world temporal networks. The description of the datasets is shown in Table I. Among them, DBLP-3 and DBLP-5***https://dblp.uni-trier.de are co-author networks that consist of authors as nodes and co-authorship as edges, with the node features being abstracts of the author’s publications during certain period encoded by word2vec. The authors in DBLP-3 and DBLP-5 are from three and five research areas, respectively. Reddithttps://www.reddit.com/ is a social network where the nodes represent the posts and edges are defined through similar keywords. Word2vec is applied to the comments of a post to generate its node attributes. Brain is a biological network, with nodes symbolizing distinct cubes of brain tissue and edges reflecting their connectivity.
@@
TABLE I: Statistics of the Datasets. AR represents the anomaly ratio, calculated as the ratio of the number of anomalies to the total number of nodes.
Datasets Nodes Edges Attributes Timestamps Anomalies
DBLP-3 4,257 38,240 100 10 210
DBLP-5 6,606 65,915 100 10 330
Reddit 8,291 292,400 20 10 420
Brain 5,000 1,975,648 20 12 240
Bitcoin-OTC 6,005 355,096 32 138 300
DGraph 3,700,550 4,300,999 17 821 15,509
Connectivity between two nodes is established if they exhibit a similar level of activation during the observed time frame. BitcoinOTC is a who-trusts-whom network of bitcoin users trading on the platforms www.bitcoinotc.com. Nodes represent the users from the platform, and an edge appears when one user rates another on the platform. DGraph [47] is a large-scale financial network, showcasing registered users as nodes, emergency contact relationships as edges, and 17 attributes from users’ personal profiles as node features.
Since only DGraph has ground-truth labels for anomalies, we manually inject synthetic anomalies into the other five datasets for evaluation in the testing phase. For a fair comparison, we follow the anomaly injection strategies used in [22] and [11], and inject equal numbers of structural anomalies and attributive anomalies for each snapshot $\mathcal{G}^{t}$ in the test set:
  • Structural anomaly injection. Following [22], we generate structural anomalies by randomly selecting $N_{p}$ nodes from node set $\mathcal{V}^{t}$ and connecting them to form fully connected cliques. The selected $N_{p}$ nodes are labeled as structural anomaly nodes. This process is repeated $q$ times to generate $q$ cliques.
  • Attributive anomaly injection. Following [11], attributive anomalies are created by randomly selecting $N_{p}\times q$ nodes from $\mathcal{V}^{t}$. For each chosen node $v^{t}_{i}$, we sample an additional $k$ nodes to form a candidate set: $\mathcal{V}^{t}_{i,attr}=\{v^{t}_{1},\ldots,v^{t}_{k}\}$. Then, we replace the feature vector of $v_{i}$ with the node feature from $\mathcal{V}^{t}_{i,attr}$ that has the largest attribute distance from $v^{t}_{i}$. Following [11], we set $k$=50 for all the datasets.
TABLE II: Node anomaly detection performance on six benchmark datasets, the best and second to best results on each dataset are in bold and underlined, respectively. PRE, F1, and AUC represent the precision, macro-F1, and area under the curve, respectively.
Dataset DBLP-3 DBLP-5 Reddit
Metrics PRE F1 AUC PRE F1 AUC PRE F1 AUC
DOMINANT 0.1358 0.5490 0.6994 0.5431 0.7327 0.9154 0.4912 0.6326 0.9316
CoLA 0.4753 0.4874 0.5814 0.4750 0.4872 0.4806 0.4747 0.4870 0.2410
SL-GAD 0.5302 0.5193 0.6174 0.5229 0.5038 0.7638 0.4908 0.4844 0.6843
GRADATE 0.5126 0.4874 0.5581 0.4923 0.4875 0.4726 0.5231 0.4870 0.5823
NetWalk 0.5336 0.5092 0.7126 0.5805 0.5034 0.9107 0.5606 0.5726 0.7821
MTHL 0.5758 0.5077 0.5901 0.5433 0.4951 0.7382 0.6295 0.6279 0.7074
TADDY 0.4994 0.5124 0.6425 0.4994 0.5135 0.6878 0.5624 0.5246 0.7390
SAD 0.7044 0.7207 0.8973 0.4931 0.3523 0.4535 0.5198 0.5219 0.8551
CLDG 0.7225 0.7188 0.8882 0.7208 0.7021 0.8781 0.6819 0.6601 0.8348
STRIPE 0.7622 0.7972 0.9620 0.7359 0.8020 0.9765 0.9409 0.6849 0.9810
Dataset Brain Bitcoin-OTC DGraph
Metrics PRE F1 AUC PRE F1 AUC PRE F1 AUC
DOMINANT 0.5845 0.5958 0.8212 0.5157 0.5062 0.9079 0.4985 0.4914 0.5709
CoLA 0.4760 0.4877 0.5716 0.4750 0.4867 0.7117 0.3312 0.3893 0.4361
SL-GAD 0.6640 0.6347 0.8735 0.4750 0.4871 0.9356 OOM OOM OOM
GRADATE 0.4910 0.4871 0.5629 0.4971 0.4868 0.7592 OOM OOM OOM
NetWalk 0.6239 0.6643 0.8590 0.6667 0.6324 0.9504 OOM OOM OOM
MTHL 0.5843 0.5988 0.8119 0.6727 0.6410 0.9353 OOM OOM OOM
TADDY 0.6218 0.6365 0.7235 0.6552 0.6821 0.9566 0.5940 0.5477 0.6654
SAD 0.5266 0.5065 0.5638 0.5733 0.7277 0.6341 0.4248 0.6136 0.7312
CLDG 0.6372 0.6425 0.5928 0.7315 0.7544 0.8394 0.6210 0.6018 0.6528
STRIPE 0.6919 0.7144 0.9389 0.7579 0.8259 0.9952 0.6451 0.6514 0.7526